OpenStudy (adamaero):

If f(x) varies directly with x2, and f(x) = 72 when x = 6, find the value of f(3) MEDAL

OpenStudy (anonymous):

ABOUT A WEEK AGO

OpenStudy (anonymous):

Varies directly just means proportional.\[ f(x)=kx^2 \]

OpenStudy (anonymous):

No^

OpenStudy (anonymous):

http://prntscr.com/5f2rg8

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

you will find what your looking for right here https://www.youtube.com/watch?v=vJwKKKd2ZYE

OpenStudy (anonymous):

TheycallmeAladdin

OpenStudy (adamaero):

do not spam

OpenStudy (anonymous):

I am not

OpenStudy (anonymous):

http://prntscr.com/5f2rg8

OpenStudy (anonymous):

http://prntscr.com/5f2rg8

OpenStudy (anonymous):

They gave us \(f(6)=72\), which means \[ 72=k(6)^2 \]Allowing us to find \(k\).

OpenStudy (anonymous):

http://prntscr.com/5f2rg8

OpenStudy (anonymous):

Or, we could have just used: \[ \frac{f(3)}{72}=\frac{f(3)}{f(6)} = \frac{(3)^2}{(6)^2} \]